OpenStudy (sonyalee77):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12?

OpenStudy (anonymous):

First you have to solve in y = mx + b format 4x-3y = -12 -3y = -4x - 12 y = 4x/3 + 4 Are you with me so far?

OpenStudy (anonymous):

If another graph is parallel to another one, they have the same slope. Any equation with the same slope will be parallel y = 4x/3 + 2 y = 4x/3 - 5